# LayerTracer: Cognitive-Aligned Layered SVG Synthesis via Diffusion Transformer

### v1

@@ 4.3. Comparison and Evaluation
Figure 5. Compare with baseline methods in Text-to-SVG generation task.
Figure 6. Compare with baseline methods in layer-wise vectorization task. Our results are more concise and exhibit more logical layering.
@@
Table 1. Comparison with SOTA SVG Generation Methods. The best results are denoted as Bold.
@@
Vecfusion (Jain et al., 2023) 31.10 4668 128.00 False
SVGDreamer (Xing et al., 2024b) 32.68 5715 512.00 False
DiffSketcher (Xing et al., 2023) 31.47 3374 512.00 False
@@
Table 2. Comparison with SOTA Vectorization Methods. The best results are denoted as Bold.
@@
Diffvg (Li et al., 2020) $2.02\times 10^{-4}$ 393 256.00 False
LIVE (Ma et al., 2022) $5.21\times 10^{-4}$ 3147 46.00 True
O&R (Hirschorn et al., 2024) $2.01\times 10^{-4}$ 612 64.00 True
@@
Figure 7. Ablation study of Serpentine Layout Strategy.
4.4. Ablation Study of Serpentine Layout Strategy.
In this section, we conduct an ablation study on the serpentine layout strategy. As shown in Fig. 7, when the serpentine layout strategy is not used to construct the training dataset, incomplete decomposition, undesirable repetitions, and abrupt changes between frames are more likely to occur. The quantitative evaluation results are presented in Table 3. For the layer-wise vectorization task, we calculate the MSE between the predicted results for 9 frames and the ground truth on both the training and test sets. When the serpentine layout strategy is not applied, the MSE is higher.
Table 3. Quantitative Evaluation of Serpentine Layout Strategy.
@@
Full $1.65\times 10^{-4}$ 0.971 $1.99\times 10^{-4}$ 0.963
4.5. User Study
@@
Figure 8. User Study Results. LayerTracer outperform in all three metrics.
5. Limitations and Future Work

### v3

@@ 4.3 Comparison and Evaluation
Figure 6: Compare with baseline methods in Text-to-SVG generation task.
Figure 7: Compare with baseline methods in layer-wise vectorization task. Our results are more concise and exhibit more logical layering.
@@
Table 1: Comparison with SOTA SVG Generation Methods. The best results are denoted as Bold.
@@
Vecfusion [18] 31.10 4668 128.00 False
SVGDreamer [58] 32.68 5715 512.00 False
DiffSketcher [57] 31.47 3374 512.00 False
@@
Table 2: Comparison with SOTA Vectorization Methods. The best results are denoted as Bold.
@@
Diffvg [20] $2.02\times 10^{-4}$ 393 256.00 False
LIVE [24] $5.21\times 10^{-4}$ 3147 46.00 True
O&R [14] $2.01\times 10^{-4}$ 612 64.00 True
@@
Figure 8: Ablation study of Serpentine Layout Strategy.
Table 3: Quantitative Evaluation of Serpentine Layout Strategy.
@@
Full $1.65\times 10^{-4}$ 0.971 $1.96\times 10^{-4}$ 0.963
4.4 Ablation Study of Serpentine Layout Strategy.
In this section, we conduct an ablation study on the serpentine layout strategy. As shown in Fig. 8, when the serpentine layout strategy is not used to construct the training dataset, incomplete decomposition, undesirable repetitions, and abrupt changes between frames are more likely to occur. The quantitative evaluation results are presented in Table 3. For the layer-wise vectorization task, we calculate the MSE between the predicted results for 9 frames and the ground truth on both the training and test sets. When the serpentine layout strategy is not applied, the MSE is higher. This indicates that the Serpentine Layout Strategy benefits the model in learning a consistent layering process.
Figure 9: User Study Results. LayerTracer outperform in all three metrics.
4.5 User Study
@@
5 Limitations and Future Work
@@ 6 Conclusion
Acknowledgement
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030).
Supplementary Material
A User Study Details
We conducted a user study in the form of an online survey, with a total of 46 participants evaluating 36 questions per questionnaire. The study was divided into two sections:
The first section evaluated text-to-vector generation results, comparing different methods. Each comparison included the corresponding text prompt, allowing participants to select the generated vector graphic that best reflected the original textual description. The evaluation consisted of three examples per method, leading to a total of nine comparisons, with two questions per comparison:
1. Which vector graphics result looks better? 2. Which result better reflects the meaning of the original text?
The second section focused on layer-wise vectorization, comparing different approaches. To ensure a fair evaluation, we provided a structured breakdown of the vectorization process, allowing participants to assess the quality and rationality of the vector graphic creation process. This section also contained three examples per method, resulting in nine comparisons, with two evaluation questions:
1. Which vectorization result looks better? 2. Which result better preserves the characteristics of the original raster image?
B Dataset Construction Methods.
The SVG dataset proposed in this work is collected from multiple sources, including the internet, vendor procurement, and designer-created assets. In addition to manually crafted layered SVGs, we also engaged designers to reorganize and arrange layers of other SVGs into a logical sequence that aligns with the design creation process. The final training data was obtained through multiple rounds of human-in-the-loop filtering. The dataset consists of a total of 20K samples across three styles: black outline icons (3K), illustrations (2K), and emojis (15K). For caption annotation, we added different triggers for different datasets and used the Florence-2 model [54] to label the last frame of the SVG sequence.
Figure 10: Examples of questions in the User Study online questionnaire.
C Failure Cases
This section presents some failure cases of LayerTracer. As shown in the figure, when generating pixel-format sequence data, issues such as unwanted repeat, no layering, and inconsistencies between consecutive frames may occur. The problem of undesired repetition can be mitigated during the Layer-wise vectorization stage by comparing adjacent frames to remove duplicates. However, incorrect layering and frame inconsistencies can negatively impact the quality of Layer-wise SVG generation.
Figure 11: More results of layered vector graphics generation. The red boxes highlight the SVG format results, which can be zoomed in to view the details.
Figure 12: Each row represents examples of SVG datasets for three categories: emojis, illustrations, and black outline icons.
Figure 13: Some failure cases.